The speed of a projectile when it is at its greatest height is (2/5)raised to the power 1/2 times its speed at half the maximum height.The angle of projection is?

Grade:12

2 Answers

Chetan Mandayam Nayakar
13 years ago

ucos(theta) = (√(2/5))(√((u2cos2(theta)+(u2sin2(theta)-(2gu2sin2(theta)/4g))))

thus, cos(theta)= sqrt((1+(cos(theta))^2)/5)

thus theta = 60 degrees

Rahul
5 years ago
V(t)=complete ✓10sqr+gt(sqr}125=100+g(sqr)× t(sqr)T(sqr=125-100/100T(sqr)= 25/100So t=0.5Thus itvis the wrie akbeosbqkby rahul mahech if any mistake in this question tell me